Stato attivita` e richieste 2014 per i rivelatori di muoni

Commenti

Transcript

Stato attivita` e richieste 2014 per i rivelatori di muoni
Stato attivita` e richieste 2014
per i rivelatori di muoni ed il trigger
D.Boscherini, G.Gaudio, A.Negri
in rappresentanza dei gruppi RPC, MDT e TDAQ
ATLAS Referees - Pisa, 05/09/2013
1 MDT
2 A%vità MDT durante LS1 • 
• 
• 
• 
• 
• 
Completamento dell'installazione, allineamento e commissioning delle camere EE side A nella regione di transizione La small wheel C è stata portata in superficie a metà marzo 2013 per consenGre le operazioni di installazione dell'IBL –  richiederà il commissioning dei rivelatori una volta che la wheel sarà riposizionata in caverna. Ordinaria manutenzione dei rivelatori, in parGcolare per l'eleLronica di front-­‐end e i sensori ambientali e di allineamento, Riparazione di perdite nel sistema di gas, in parGcolare per le camere EO (crack dei jumpers), e per altri pochi casi isolaG –  importante contributo italiano per l'experGse accumulata Installazione e commissioning delle camere negli ascensori, MDT e sMDT + RPC; –  possibile contributo da esperG italiani E' stato osservato un problema di saturazione ad alta luminosità istantanea sulle camere MDT interne della small wheel: è in corso uno studio per risolvere il problema A%vità MDT durante LS1 •  Aggiornamento DCS (hardware e soZware): –  entro il 2013 sosGtuzione tu% i mainframe Caen SY1527 con la nuova versione SY4527 e tuLe le ventole di (praGcamente) tuLe le alimentazioni dei crate Wiener. –  Queste a%vità rientrano nel programma di consolidamento e mantenimento del sistema di muoni e quindi saranno pagate sui fondi M&OB. Da tarda primavera 2014: •  recommissioning di tuLo il sistema –  Rivelatori –  Acquisizione daG ( responsabilità italiana) –  Data quality, sia online che offline. •  ripresa delle a%vità ai calibraGon center per gli MDT (Roma) –  procedura di calibrazione e data quality –  o%mizzazione del soZware per il calcolo e l'uGlizzo delle costanG, –  migrazione dei database oracle di calibrazione nei vari siG 5 RPC
Reminder: RPC in ATLAS sono 100% italiani, quindi tu;a l’a=vita` riportata e` italiana 6 Status of ongoing activities /1
Gas leak repairs
New repair techniques developed and already used (see next slide)
Start-up time spent to develop and optimize the techniques
Repair of known leaks in progress:
-  6/60 BOL leaks repaired (left behind because of difficult access, now
intervention possible with the high-quality endoscope provided by ATLAS TC)
Gas leak search
Gas group flowmeters providing very rough measurements (known leaks
accounting for only half of the total leak rate)
Complete check-up of the full system on-going:
-  32/128 manifolds measured using a manometer on each isolated manifold
à  producing a list of (up-to-now) unknown leaks
7 Gas leak repair techniques
Incomplete crack complete breakdown spray glue Install new inlet 8 Status of ongoing activities /2
Double chambers in toroid support feet
Cable installation completed
(a few flat cables, HV connectors,
power cables were missing)
-  HV- and LV-splitters ordered
-  new channel map under construction
-  cable testing needed
additional RPC chambers in BO stations
32/32 new trigger boxes installed
32/32 new splitter boxes installed
9 Status of ongoing activities /2
Chambers in the elevator regions
BME
BOE
BME
BOE
BOE to be installed in next October + November
Gas volumes with 2mm gap, the same as the rest of ATLAS
Chambers under construction at General Tecnica srl,
expected delivery: next week
BME installation currently planned in next March
New generation gas volumes with 1mm gaps will be used
Design is starting
10 Status of ongoing activities /3
Noise reduction
Enhance of Faraday cage in chambers with panels operated at harder Vth
due to e.m. pickup noise (~200 panels out of 3000)
Installation (and gluing) of cable-stops
in front-end output cables is just started (1 sector)
Expected to finish end 2013 – beginning 2014
DCS upgrade
Adjusting the DCS to the LHC high luminosity period will require:
-  test with new CAEN mainframe done
-  new hardware for double chambers included
-  map for new channels under construction
Still to be done (2013-2014):
-  replacement of obsolete PCs
-  implementation of luminosity based alarms and settings
-  integration of LVL1/pad info into the DCS
-  enhancement of automatic alarms/monitoring/analysis
-  improvement of RPC/common muon infrastructure (rack reset network,
service voltage monitoring)
11 Status of ongoing activities /4
Trigger upgrade Usage of the new CAN interfaces (Peak/Systec) on SLC6 was successful and now moving to configuraGon s/w To be done (2013-­‐2014): -­‐  Rewrite configuraGon soZware -­‐  New auto-­‐recovery procedures in case of LVL1 misalignment, and in general more robustness in case of hardware failure 12 RPC:
plans and requests for 2014
13 Upgrade of gas flowmeters
Current setup:
-  128 input manifolds with one flowmeter each
-  128 output manifold with one flowmeter each
Each manifold serves one RPC layer to 6 (BM) or to 12 (BO) chambers
i.e. gas flow measurements integrated over many gas volumes
input layer 1 1 gas line / layer in BM staGons output layer 0 side C side A output layer 1 input layer 0 input layer 1 output layer 0 2 gas lines / layer in BO staGons input layer 1 output layer 0 side C output layer 1 side A output layer 1 14 Upgrade of gas flowmeters
Proposal:
-  install one flowmeter on each chamber output (one per layer)
~1200 flowmeters in total
Advantages:
-  spot leak in its early stage, i.e. easier repair
-  reduce leak rate
-  reduce time spent in cavern
-  improve gas flow monitoring, help in understanding gap problems
Request:
-  40 kE flowmeters
-  20 kE readout with ADC boards
-  20 kE cables
Total request: 80 kE
15 Activities and requests for 2014 /1
Gas repairs Only half of the total leak rate is due to know leaks Finding and repair of leaks over the 128 manifolds 4m x 2 persons à 8m FTE PrevenGon work to safely operate the full gas system 2m x 2 persons à 4m FTE Total = 12m FTE Replacement of gas impedances Current impedances unsuitable for high gas flow (linearity loss) and for high luminosity (need to increase the weight of background wrt chamber volume). Replacement of 2000 impedances to be scheduled 1m x 2 persons à 2m FTE Upgrade of gas flowmeters Foreseen acGvity: 16sect x 1w x 2FTE à 6m FTE Planning to get 1m FTE from M&O à 5m FTE 16 Activities and requests for 2014 /2
Upgrade in the toroid support feet Commissioning compleGon of the addiGonal 20 RPC chambers i.e. debugging of chambers and tools needed (DCS, DAQ, DQ) 3m FTE Upgrade in the elevator regions InstallaGon of BME chambers (BOEs will be installed in 2013) and commissioning of the whole project: -­‐  integraGon with MDT -­‐  cabling and power-­‐on -­‐  cosmic rays test -­‐  final debugging 3m FTE (2m FTE received in 2013) 17 Activities and requests for 2014 /3
DCS upgrade CompleGon of the upgrade to high luminosity and integraGon of the LVL1 info into DCS (only trigger rates from low-­‐pt towers in 2012) 3m FTE Data-­‐taking restart Detector + DCS + DAQ + DQ 2mu x 4FTE à 8m FTE (+ 2-­‐4 weeks of data-­‐taking shiZs not considered) CoordinaWon CoordinaGon of technicians by acGvity responsibles: Boscherini(gen.), Sekhniaidze(gas), Cardarelli(FEE/power-­‐system), Aielli(DCS) Total à 4m FTE 18 Richiesta per upgrade flussimetri
Item kE Flowmeters 40 ADC boards 20 Cables 20 Total 80 19 Richiesta missioni 2014
AcWvity FTE (months) Gas leak repairs 12 Replacement of gas impedances 2 Gas flowmeter upgrade 5 Upgrade feet regions 3 Upgrade elevator regions 3 DCS upgrade 3 Data taking restart 8 AcGvity coordinaGon 4 Total 40 20 Responsabilita`
IsWtuto BO LE NA RM2 Nome A=vita` Livello D.Boscherini Coordinatore RPC + RPC responsible in the Muon Steering Group L2 D.Boscherini Maintenance overview L3 A.Bruni DQ online presenter and automaGc checks L3 G.Chiodini Offline data quality monitor and assessment L3 G.Chiodini/S.Spagnolo RPC simulaGon L3 S.Spagnolo RPC soZware descripGon L3 M.Della Pietra Data quality coordinator L3 M.Della Pietra Gnam (online monitoring) L3 G.Sekhniaidze Gas system coordinator L3 G.Aielli DCS coordinator L3 R.Cardarelli Front-­‐end electronics and power system L3 R.Santonico Gas repair task force coordinator L3 21 TDAQ
22 T/DAQ: stato e richieste 2014 •  InfrastuLura network @P1 rinnovata •  DAQ: accorpamento farm di L2 & EB & EF [Coordinazione PV] -­‐  Stato dell'a%vità in linea con i programmi -­‐  Nuovo design validato con successo in technical run dedicaG @P1 [Responsabilità RM1] -­‐  AdaLamento dell'infrastruLura di muon calibraGon stream in corso •  HLT: [LE, RM1, GE con molti ruoli di coordinamento] -­‐  Aggiornamento infrastruLura per accorpamento L2&EF a buon punto •  Comporta cambiamenG significaGvi in quasi tuLe le “segnature” Studi sull'uGlizzo delle informazioni di FTK in corso -­‐ 
-­‐  Si sta lavorando per migliorare velocità e prestazioni degli algoritmi -­‐  In parallelo, proseguono gli studi di performance ed efficienza sui daG 2012 (es: muon, b-­‐tagging, ..) e la preparazione delle relaGve note •  Nessuna richiesta per il 2014 •  Riconosciuto 1 FTE di contributo in-­‐kind per sys-­‐admin (92 kCHF) 26/08/13 Andrea Negri @Atlas-­‐It 23 L1 mu-­‐barrel: stato e richieste 2014 •  Consolidamento regione piedi spettrometro [RM1] -  elettronica on-detector installata, cabling da finire entro l'anno -  Richieste: 4 MU •  2 per test elettronica on-detector, 2 per calibrazione e commissioning •  Upgrade regione ascensore [RM1] -  Installazione elettronica, validazione e integrazione nel trigger -  Richieste: 4 MU •  2 per test elettronica on-detector, 2 per calibrazione e commissioning •  Upgrade HW inizializzazione dell'elettronica [RM1] -  nuova elettronica arrivata, da installare entro fine anno -  Richieste: 1MU per installazione nuovo sistema •  Upgrade SW/DAQ: [BO, NA, RM1, RM2] -  Attività avviata, prima versione entro l'anno, commissioning nel 2014 -  Richieste: 1MU per test nuovo SW e integrazione nel DAQ •  Commissioning L1 muon barrel per preparazione presa dati run2 [RM1] -  Richieste: 1 MU 26/08/13 Andrea Negri @Atlas-­‐It 24 T/DAQ: responsabilità 26/08/13 Andrea Negri @Atlas-­‐It 25 

Documenti analoghi